# MaIR: A Locality- and Continuity-Preserving Mamba for Image Restoration

### v1

@@ Abstract
Recent advancements in Mamba have shown promising results in image restoration. These methods typically flatten 2D images into multiple distinct 1D sequences along rows and columns, process each sequence independently using selective scan operation, and recombine them to form the outputs. However, such a paradigm overlooks two vital aspects: i) the local relationships and spatial continuity inherent in natural images, and ii) the discrepancies among sequences unfolded through totally different ways. To overcome the drawbacks, we explore two problems in Mamba-based restoration methods: i) how to design a scanning strategy preserving both locality and continuity while facilitating restoration, and ii) how to aggregate the distinct sequences unfolded in totally different ways. To address these problems, we propose a novel Mamba-based Image Restoration model (MaIR), which consists of Nested S-shaped Scanning strategy (NSS) and Sequence Shuffle Attention block (SSA). Specifically, NSS preserves locality and continuity of the input images through the stripe-based scanning region and the S-shaped scanning path, respectively. SSA aggregates sequences through calculating attention weights within the corresponding channels of different sequences. Thanks to NSS and SSA, MaIR surpasses 40 baselines across 14 challenging datasets, achieving state-of-the-art performance on the tasks of image super-resolution, denoising, deblurring and dehazing. Our codes will be available after acceptance.
@@ 3.2 MaIR Module
SSA: SSA aggregates the processed sequences by calculating attentions within corresponding channels. This design enables it to capture complex dependencies across distinct sequences, thus better leveraging complementary information from different scanning directions.
As shown in Fig. 4, supposing sequence number $K=4$, for SSO-processed sequences $\{X^{i}\}_{i=1}^{4}$, we first apply spatial average pooling $\Phi_{AP}(\cdot)$ to reduce the computational cost, and then concatenate as
@@ 4.5 Analysis Experiments
In this section, we first conduct ablation studies to verify the effectiveness of the NSS and SSA. Then, we introduce analysis experiments to investigate the impact of stripe width on the overall performance.
Table 7: Ablation study on the proposed NSS scheme, tested on lightweight super-resolution tasks with scale factor ×$\times 2$. The results on the Urban100 dataset are presented, which demonstrates the effectiveness of the NSS.
Baseline w/o NSS w/o SS LM ZigMa Hilbert
@@ 4.5.1 Ablation Studies
We first conduct ablation study to analyze the effectiveness of NSS. In detail, five configurations are conducted, i) replacing NSS with Z-shaped scanning strategy (denoted as w/o NSS), ii) removing shift stripe (denoted as w/o SS), iii) replacing NSS with the scanning strategy in LocalMamba [22] (denoted as LM), iv) replacing NSS by the scanning strategy in ZigMa [20] (denoted as ZigMa) and v) replacing NSS by the Hilbert curve (denoted as Hilbert). As illustrated in Tab. 7, one could see that NSS is important to improve the performance of MaIR.
Table 8: Ablation study on SSA for lightweight super-resolution with scale factor ×$\times 2$. The results on Urban100 dataset are presented, which demonstrates the effectiveness of the NSS.
Baseline w.o. SSA Gating
PSNR 32.97 32.90 32.92
SSIM 0.9359 0.9350 0.9252
To investigate the effectiveness of SSA, we also conduct experiments under the following settings: i) removing proposed SSA and aggregating sequences via sequences-wise addition (termed as w.o. SSA), and ii) replacing SSA by the gating strategy in [5] (termed as Gating). As reported in Tab. 8, SSA is more effective than others.
@@ 4.5.2 Results on different stripe width
To investigate influence of stripe width, we train lightweight SR model with various stripe width (2, 4, 8, 16 and 32) and evaluate them on Urban100 dataset. As presented in Tab. 9, one could observe that the PSNR and SSIM values are quite similar under different settings, except for the cases with the largest and the smallest stripe widths. It indicates that the proposed method exhibits robustness against changes in stripe width, maintaining high-quality image restoration across a range of stripe widths.
@@ 5 Conclusion
In this paper, we propose MaIR, a novel state space model for image restoration that can preserve both local dependencies and spatial continuity of input images. To this end, we propose two designs: Nested S-shaped Scanning strategy (NSS) and Sequences Shuffle Attention (SSA). NSS is designed to extract locality- and continuity-preserving sequences from images, and SSA adaptively aggregates these sequences. Thanks to their cooperation, MaIR not only addresses the limitations of existing Mamba-based restoration methods but also improves image quality without introducing extra computations. Extensive experiments across four tasks on 14 benchmarks comparing with 40 baselines, validate the superiority of MaIR, demonstrating its robustness and effectiveness in various image restoration tasks.

### v2

@@ Abstract
Recent advancements in Mamba have shown promising results in image restoration. These methods typically flatten 2D images into multiple distinct 1D sequences along rows and columns, process each sequence independently using selective scan operation, and recombine them to form the outputs. However, such a paradigm overlooks two vital aspects: i) the local relationships and spatial continuity inherent in natural images, and ii) the discrepancies among sequences unfolded through totally different ways. To overcome the drawbacks, we explore two problems in Mamba-based restoration methods: i) how to design a scanning strategy preserving both locality and continuity while facilitating restoration, and ii) how to aggregate the distinct sequences unfolded in totally different ways. To address these problems, we propose a novel Mamba-based Image Restoration model (MaIR), which consists of Nested S-shaped Scanning strategy (NSS) and Sequence Shuffle Attention block (SSA). Specifically, NSS preserves locality and continuity of the input images through the stripe-based scanning region and the S-shaped scanning path, respectively. SSA aggregates sequences through calculating attention weights within the corresponding channels of different sequences. Thanks to NSS and SSA, MaIR surpasses 40 baselines across 14 challenging datasets, achieving state-of-the-art performance on the tasks of image super-resolution, denoising, deblurring and dehazing. The code is available at https://github.com/XLearning-SCU/2025-CVPR-MaIR.
@@ 3.2 MaIR Module
SSA: SSA aggregates the processed sequences by calculating attentions within corresponding channels. This design enables it to capture complex dependencies across distinct sequences, thus better leveraging complementary information from different scanning directions. As shown in Fig. 4, supposing sequence number $K=4$, for SSO-processed sequences $\{X^{i}\}_{i=1}^{4}$, we first apply spatial average pooling $\Phi_{AP}(\cdot)$ to reduce the computational cost, and then concatenate as
@@ 4.5 Analysis Experiments
In this section, we first conduct ablation studies to verify the effectiveness of the NSS and SSA. Then, we introduce analysis experiments to verify the observations and investigate the impact of stripe width on the overall performance.
Table 7: Ablation study on NSS, tested on lightweight super-resolution with scale factor ×$\times 2$. The results on the Urban100 dataset are presented, which demonstrates the effectiveness of the NSS.
Baseline w/o NSS w/o SS LM ZigMa PH
@@ 4.5.1 Ablation Studies
We first conduct ablation study to analyze the effectiveness of NSS. In detail, five configurations are conducted, i) replacing NSS with Z-shaped scanning strategy (denoted as w/o NSS), ii) removing shift stripe (denoted as w/o SS), iii) replacing NSS with the scanning strategy in LocalMamba [21] (denoted as LM), iv) replacing NSS by the scanning strategy in ZigMa [19] (denoted as ZigMa) and v) replacing NSS by the Peano-Hilbert curve (denoted as PH). As illustrated in Tab. 7, NSS is important to improve the performance of MaIR.
Table 8: Ablation study on SSA for lightweight super-resolution with scale factor ×$\times 2$. The results on Urban100 dataset demonstrate the effectiveness of SSA.
MaIR w/o SSA UVM SeqGat
PSNR 32.97 32.90 32.59 32.92
SSIM 0.9359 0.9350 0.9324 0.9352
CAGat FPixGat DWPixGat -
PSNR 32.71 31.90 32.74 -
SSIM 0.9335 0.9250 0.9342 -
To investigate the effectiveness of SSA, we remove SSA and aggregate sequences through: i) sequences-wise addition (termed as w/o SSA), ii) SSM [67] (termed as UVM), iii) sequence-wise gating [5] (termed as SeqGat), iv) channel-wise gating (termed as CAGat), v) pixel-wise gating through fully connected convolution (termed as FPixGat). vi) pixel-wise gating through depth-wise convolution (termed as DWPixGat). It is worth noting that we keep the size of different models to be similar for fair comparisons. As shown in Tab. 8, SSA is more effective than others.
@@
Figure 9: Visual comparisons of different scanning strategies, illustrating that i) windows-based scanning path overlooks the continuity between different regions (e.g., relationship between different layers of the scarf), resulting in wrong textures, ii) S-shaped scanning path leads to distortion in local regions, causing the scarf’s texture to appear warped. iii) Z-shaped scanning path suffers from both of them. In contrast, MaIR avoids aforementioned problems and achieves visually appealing results.
4.5.2 Verification of Observations
To verify observations shown in Fig. 1, we conduct visual comparisons among different scanning strategies. As shown in Fig. 9, proposed methods can maintain both locality and continuity and produce more visual pleasant results.
@@ 4.5.3 Results on different stripe width
To investigate influence of stripe width, we train lightweight SR model with stripe width $w_{s}=\{2,4,8,16,32\}$ and evaluate them on Urban100 dataset. As presented in Tab. 9, the PSNR and SSIM values are quite similar under different settings, except for the cases with the largest and the smallest stripe widths. It indicates that the proposed method exhibits robustness against changes in stripe width, maintaining high-quality image restoration across a range of stripe widths.
@@ 5 Conclusion
In this paper, we propose MaIR, a novel state space model for image restoration that can preserve both local dependencies and spatial continuity of input images. To this end, we propose two designs: Nested S-shaped Scanning strategy (NSS) and Sequences Shuffle Attention (SSA). NSS is designed to extract locality- and continuity-preserving sequences from images, and SSA adaptively aggregates these sequences. Thanks to their cooperation, MaIR not only addresses the limitations of existing Mamba-based restoration methods but also improves image quality without introducing extra computations. Extensive experiments across four tasks on 14 benchmarks comparing with 40 baselines validate the superiority of MaIR, demonstrating its robustness and effectiveness in various image restoration tasks.
Acknowledgments
This work was supported in part by NSFC under Grant 62176171, U21B2040, 62472295; in part by the Fundamental Research Funds for the Central Universities under Grant CJ202303; and in part by Sichuan Science and Technology Planning Project under Grant 24NSFTD0130.